OpenStudy (anonymous):

3^2+2*2^2 169 25 44 17

OpenStudy (anonymous):

17

OpenStudy (anonymous):

What is 3^2? What is 2^2?

OpenStudy (anonymous):

Roachie - don't just give answers please.

OpenStudy (anonymous):

8+9=17

OpenStudy (anonymous):

3^2+2*2^2 =9+2*4=9+8=17